OpenStudy (anonymous):

subtract 4/x+1 - x+3/x

OpenStudy (anonymous):

\[\frac{4}{x+1}-\frac{x+3}{x}\]

OpenStudy (anonymous):

that one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{4x-(x+1)(x-3)}{(x+1)x}\]

OpenStudy (anonymous):

then some algebra

OpenStudy (anonymous):

is it x^2 +3/ x^2 +x

OpenStudy (anonymous):

numerator is \[4x-(x^2-2x-3)\] \[-x^2+6x+3\] denominator is \[x^2+x\]

OpenStudy (anonymous):

for a 'final answer" of \[\frac{-x^2+6x+3}{x^2+x}\]

OpenStudy (anonymous):

4x+1−x+3x

OpenStudy (anonymous):

dont forget medal